### v2

# Pinching-Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

Abstract
This article proposes a novel design for the Pinching Antenna SyStems (PASS) and advocates simple yet efficient wireless communications over the “last meter”. First, the potential benefits of PASS are discussed in the paper by reviewing an existing prototype. Then, the fundamentals of PASS are introduced, including their physical principles, signal models, and communication designs. In contrast to existing multi-antenna systems, PASS brings a novel concept, termed Pinching Beamforming, which is achieved by dynamically adjusting the positions of pinching antennas (PAs). Based on this concept, a couple of practical transmission architectures are proposed for realizing the full potentia of PASS, namely non-multiplexing and multiplexing architectures. More particularly, 1) the non-multiplexing architecture is featured by simple baseband signal processing and relies on the pinching beamforming only; while 2) the multiplexing architecture provides enhanced signal processing capabilities with joint baseband and pinching beamforming, which can be further divided into sub-connected, fully-connected, and phase-shifter-based fully-connected schemes. Furthermore, several emerging scenarios are put forward for integrating PASS into future wireless networks. As a further advance, by demonstrating a few numerical case studies, the significant performance gain of PASS is revealed compared to conventional multi-antenna systems. Finally, several research opportunities and open problems of PASS are highlighted.
I Introduction
The perspective introduced by Shannon’s information theory established the foundation for the digital revolution and has profoundly shaped the development of modern communication systems [1]. In Shannon’s original framework, the channel connecting the transmitter and receiver was treated as a fixed entity, i.e., a parameter entirely dictated by environmental factors and beyond human control. However, advances in communication technology have challenged this assumption by enabling active reconfiguration of channel parameters to boost the channel throughput. A notable example is the deployment of multiple-input multiple-output (MIMO) systems, which reconfigures the channel by creating separated parallel point-to-point links to enhance communication efficiency [2].
Building on this momentum, recent advancements in intelligent and flexible antenna designs have further introduced novel approaches for channel reconfiguration. Technologies, such as reconfigurable intelligent surfaces [3, 4], dynamic metasurface antennas [5], fluid antennas [6], and movable antennas [7], can dynamically adjust their positions, apertures, and electromagnetic properties, enabling precise control of wireless channel conditions. By optimizing their reconfigurable parameters, these novel antenna types and array architectures offer flexibility in controlling the effective end-to-end channel gain. This capability has demonstrated great potential for improving system performance and has attracted considerable research interests.
Despite their potentials to enhance network throughput, existing flexible-antenna systems face certain limitations. Firstly, those existing antenna reconfigurations are often confined to apertures spanning only a few wavelengths, which limits their ability to mitigate large-scale path loss. Secondly, in existing reconfigurable systems, it is costly, if not impossible, to change their antenna configuration, i.e., add or remove antennas once these systems are established. To address these challenges, a more flexible yet efficient new design—Pinching-Antenna SyStems (PASS) come into researchers’ sight [8, 9].
Figure 1: Comparison between PASS and conventional MIMO systems, along with an illustration of the physical principles underlying PAs and NTT DOCOMO’s prototype (Photo credit: NTT DOCOMO) © CCBY [8].
I-A The Road to Near-Wired: A Prototype of PASS
The first prototype of PASS was developed and demonstrated by NTT DOCOMO in 2021 [8], as illustrated in Figure 1. A PASS consists of a specialized transmission line known as a dielectric waveguide. Similar to leaky-wave antennas, electromagnetic waves within the waveguide are emitted by pinching small separate dielectric particles at specific points along its length. These dielectrics are typically attached to the tips of plastic clips or clothespins, forming structures known as pinching antennas (PAs). By leveraging the waveguide’s ability to transmit or receive radio waves at any pinched point, PASS enables wireless communication precisely where needed [8].
The waveguide can be pinched at multiple points to provide coverage in specific locations, allowing for flexible and dynamic antenna deployment. Signal emission at a particular location ceases immediately upon removing the pinching antenna, making the process as simple as adding or removing a clothespin from a clothesline, as illustrated in Figure 1. This adaptability enables PASS to establish communication zones tailored to users’ locations and environmental conditions, offering exceptional flexibility in antenna deployment [8].
Unlike conventional flexible-antenna systems, the length of a PASS waveguide can be extended arbitrarily, allowing PAs to be positioned closer to users to establish strong line-of-sight (LoS) links [9]. This adaptability effectively creates a near-wired connection, ensuring favorable channel conditions and shifting the focus of wireless communication networks from the “last mile” to the more manageable “last meter”. This unique capability is reflected in the name “PASS”, highlighting its ability to provide the shortest and most efficient wireless connections between the PAs and the wireless transceivers, which effectively pass wireless signals with minimal loss.
I-B Key Advantages for Employing PASS in Wireless Communications
Considering the above unique features, the key advantages of PASS are summarized as follows:
  • Enhanced LoS Stability and “Last Meter” Communications: By extending the waveguide and enabling the flexible deployment of PAs, PASS effectively mitigate large-scale path loss and establish stable LoS communication environments. This capability enhances “last meter” communications and expands network coverage, which is not achievable with other flexible-antenna technologies.
  • Flexible Pinching Beamforming Gain: PASS control signals by optimizing the locations and the numbers of PAs, which enables highly flexible and scalable pinching beamforming. This capability allows for precise management of signal strength and interference, effectively meeting stringent service requirements.
  • Utilization of Near-Field Benefits: The extended aperture of the waveguide places users within the near-field region of PASS, which enables the use of near-field electromagnetic properties, such as beam focusing, to harvest the spatial degrees of freedom and enhance wireless transmission efficiency [10].
  • Cost-Effective and Scalable Implementation: PASS are simple to implement and economical to deploy, not only because of the low cost of waveguides, but also because PASS only require adding or removing dielectric materials at selected points along the waveguide [8], which makes PASS a scalable choice for practical network design.
These advantages have fueled growing research interest in PASS. The performance of PASS has been analyzed in terms of its ergodic rate [9] and its array gain [11]. Furthermore, various activation algorithms have been proposed to optimize the locations of PAs along a single waveguide; see [9] and relevant references. However, despite these initial efforts, the design of PASS for wireless communications remains in its early stages. This gap motivates us to provide a systematic introduction to PASS. In this article, we first discuss how PASS differ from conventional multi-antenna systems. We then propose several novel beamforming architectures to facilitate the design of pinching beamforming. This is followed by an exploration of promising application scenarios for PASS and an evaluation of their performance. Finally, we highlight the open challenges of PASS to inspire future research in this promising research area.
II Pinching Antenna Fundamentals:
What Will be Different?
In this section, we introduce the fundamentals of PAs and highlight the basic differences from the conventional electronic antennas from their physical principle, signal model, and communication system design perspective, respectively.
II-A Physical Principle
Unlike conventional electronic antennas which rely on electronic devices to convert alternating electric currents into electromagnetic (EM) waves, the physical principle of PAs is based on the phenomenon of EM coupling that arises when a separate dielectric is brought into close proximity with a dielectric waveguide [8]. This coupling enables the power exchange between the waveguide and the separate dielectric, as illustrated in the bottom-right of Fig. 1. Therefore, by pinching a separate dielectric onto the waveguide, the EM waves originally confined within the waveguide are coupled into the separate dielectric, where they are subsequently radiated into free space, thereby forming an “antenna”. Unlike conventional electronic antennas, PAs are entirely passive, simpler to manufacture and deploy, and hence provide greater position flexibility.
By modeling the pinching antenna as an open-ended directional coupler, the power exchange caused by the EM coupling can be characterized by the coupled-mode theory. Let $P_{\mathrm{guide}}$ and $P_{\mathrm{pinch}}$ denote the normalized power in the waveguide and the pinched separate dielectric, respectively. Their relationship is approximately given by [12]
$P_{\mathrm{guide}}$ $=1-F\sin^{2}(\kappa L),\quad P_{\mathrm{pinch}}=F\sin^{2}(\kappa L),$
where $L$ is the coupling length shown in Fig. 1, $\kappa$ is the coupling coefficient, and $F\leq 1$ is the maximum coupling efficiency. The factors $\kappa$ and $F$ are influenced by various factors, including the effective refractive index and the transverse modes of both the waveguide and the separate dielectric. It can be observed from the above formulas that the power emitted by a PA can be controlled by adjusting the coupling length $L$. When the waveguide and the separate dielectric share the same effective refractive index, the coupling efficiency $F$ reaches its maximum value of $1$, making it possible to radiate full power from a single PA with a coupling length of $\pi/(2\kappa)$.
II-B Signal Model
As shown in Fig. 1, the composite channel from the base station (BS) to the user in the PASS system consists of two components: in-waveguide propagation and free-space propagation. Compared to free-space propagation, in-waveguide propagation incurs almost negligible path loss, e.g., approximately 0.01-0.03 dB/m for a circular copper waveguide at 15 GHz [13]. Consequently, when multiple PAs are deployed on a waveguide, the signal received at the user from the $n$-th PA through the LoS channel can be expressed mathematically as:
$y_{n}=\frac{\beta_{n}\sqrt{P_{n}}}{r_{n}}e^{-j\frac{2\pi}{\lambda}\left(r_{n}+n_{\mathrm{eff}}d_{n}\right)}x.$
In this expression, $x$ is the signal fed into the waveguide, $\beta_{n}$ is the complex channel of the free-space propagation at the reference distance of 1 meter, $P_{n}$ is the power radiated from the PA, $\lambda$ denotes the signal wavelength, $n_{\mathrm{eff}}$ represents the effective refractive index of the waveguide, $r_{n}$ is the distance of the free-space propagation from the PA to the user, and $d_{n}$ is the distance of the waveguide propagation. By adjusting the location of PAs, both $r_{n}$ and $d_{n}$ can be modified, which in turn affects the amplitude and phase of the received signal.
Based on the physical principle of PAs, the power model of PASS, signified by $P_{n}$, is significantly different from that of conventional multi-antenna systems. Specifically, with conventional antennas, the power of the signal emitted from each antenna is actively controlled by electronic devices including radio-frequency (RF) chains and power splitters. In contrast, for PAs, the transmit power is passively determined by the coupling length with the waveguide. Furthermore, when multiple PAs are pinched along a waveguide, the power radiated by subsequent antennas is influenced by the power radiated by preceding ones, leading to a coupled power model. While the coupling length of a PA is not as easily adjustable as the electronic parameters of conventional antennas since PAs are typically passive, it is still possible to design and manufacture PAs with specific coupling lengths to achieve a desired power model. In the following, we introduce two common power models for PAs.
  • Equal Power Model: In this model, the PAs on the waveguide radiates an equal amount of power. This can be achieved by designing the PAs with varying coupling lengths, such that the earlier PAs radiate a smaller proportion of the power in the waveguide, while the later PAs radiate a larger proportion of the remaining power.
  • Proportional Power Model: In this model, each PA is designed with an identical coupling length, thereby radiating the same fixed proportion of the power left in the waveguide. As a result, the power radiated by each PA decreases progressively along the waveguide, where the power radiated by later PAs is proportional to that of earlier ones by a fixed ratio.
These two power models offer different pros and cons. Compared to the proportional power model, where the latter PAs may radiate only a smaller amount of power, the equal power model ensures that all PAs are effectively utilized. The equal power model also simplifies the performance analysis of the PASS system due to its simple mathematical formulation, which is useful to identify the fundamental limits of PASS, and hence has contributed to its widespread adoption in the literature. In contrast, the proportional power model requires less stringent manufacturing specifications, as all PAs share the same design, thus reducing production complexity and costs.
Figure 2: Non-Multiplexing transmission architectures for PASS
II-C Communication System Design
From a communication system design perspective, it is crucial to position the PAs at suitable locations to maximize communication performance. On the one hand, the PAs should be deployed close to the user to mitigate the significant signal attenuation caused by free-space path loss and blockage. On the other hand, when multiple PAs are deployed along the waveguide, they form an antenna array, enabling pinching beamforming to enhance signal strength and reduce interference. By fine-tuning the positions of PAs along the waveguide, the phase shifts of the signals, due to the in-waveguide and free-space propagation, can be adjusted, thereby optimizing the beamforming gain. Leveraging the flexibility and portability of PAs, their positions can be determined either actively by the BS or passively by the user, leading to the following two designs:
  • BS-centric Design: In this design, the positions of PAs are determined by the BS to optimize specific performance metrics of the overall PASS system, such as spectral efficiency and user fairness. In dynamic scenarios, the positions of PAs can be adjusted actively based on the locations of all users in two ways: either by moving them along a pre-installed track parallel to the waveguide for continuous activation, or by selectively activating PAs at the pre-configured locations on the waveguide for discrete activation.
  • User-centric Design: In this design, users can pinch or release portable PAs onto the waveguide as needed to enhance communication service in their local area. Whenever a PA is pinched or released by a user, the baseband processing strategy needs to be optimized and updated based on the new propagation environment to avoid the performance degradation to other users.
By comparing these two designs, it can be concluded that the BS-centric design offers high flexibility, allowing the PASS to effectively mitigate path loss and blockage issues while enhancing beamforming gain for all users. However, this comes with a higher implementation cost. In contrast, the user-centric design has a lower implementation cost but may result in each PA being effectively utilized for only a single user.
III Transmission Architectures for PASS
In this section, we discuss the viable transmission architectures for PASS, while identifying their respective advantages and disadvantages.
III-A Non-Multiplexing Architecture
A schematic illustration of the non-multiplexing architecture is shown in Fig. 2. In this architecture, the data streams to be transmitted are not multiplexed at the baseband, meaning that each waveguide is fed individually by a single data stream. In this setup, the primary role of the baseband processing is to manage the power allocation across these data streams to optimize transmission efficiency. The key benefit of this design is to reduces implementation complexity. Compared to the conventional multi-antenna systems, the unique capability of pinching beamforming, achieved by adjusting the positions of PAs, can be utilized to enhance communication performance. This non-multiplexing architecture is particularly well-suited for those commonly encountered scenarios, where each waveguide serves geographically isolated areas, minimizing the need for advanced interference management between waveguides.
Figure 3: Multiplexing transmission architectures for PASS
Figure 4: Outdoor/Indoor deployments and emerging applications of PASS in 6G and beyond.
III-B Multiplexing Architecture
As shown at the bottom of Fig. 3, the multiplexing architecture multiplexes multiple data streams before feeding them into the waveguides, thus facilitating the joint transmit and pinching beamforming. This architecture can be seen as an extension of the conventional MIMO transmission architecture, where each electronic antenna is replaced by a waveguide. The key difference between the two architectures lies in the point of signal radiation. In conventional MIMO systems, signals are radiated directly from the electronic antennas into free space. In contrast, in the multiplexing architecture of PASS, signals are radiated at specific locations along the waveguide via PAs. Consequently, unlike conventional MIMO, which relies on a massive number of electronic antennas to achieve high spatial degrees of freedom (DoFs) [14], the multiplexing architecture of PASS significantly reduces the requirement for numerous waveguides because a large number of low-cost PAs can be easily installed on a single waveguide. Based on the above discussion, the multiplexing architecture for PASS can be designed in analogy to conventional MIMO architecture but with distinct differences in implementation and design focus. In the following, we introduce three basic multiplexing architectures, as shown in Fig. 3, namely sub-connected architecture, fully-connected architecture, and phase shifter (PS)-based fully-connected architecture.
  • Sub-connected Architecture: This is the simplest multiplexing architecture, where each waveguide is connected to a dedicated RF chain and carries signals that are multiplexed and precoded by baseband digital beamforming. While this architecture is analogous to the fully-digital architecture used in conventional MIMO systems, it is referred to as “sub-connected” because the signals from each RF chain are emitted exclusively by those PAs deployed on the corresponding waveguide.
  • Fully-connected Architecture: In this architecture, the signals from each RF chain are fed into all waveguides with the aid of power splitters. The fully-connected architecture addresses two key limitations of the sub-connected design. First, it enables each RF chain to connect to all PAs across all waveguides, thereby fully using the overall spatial DoFs to enhance communication performance. Second, the number of RF chains no longer needs to match the number of waveguides. This flexibility allows for either reducing the number of RF chains to save energy, as hybrid beamforing in millimeter-wave networks, or increasing it to further exploit the additional spatial DoFs provided by the massive PAs.
  • PS-based Fully-connected Architecture: This architecture builds on the fully-connected design by introducing phase shifters (PSs) to configure the signals before feeding them into the waveguides. This enhancement enables the realization of tri-hybrid beamforming, which integrates baseband digital beamforming, analog beamforming enabled by the PSs, and pinching beamforming facilitated by the PAs, exhibiting the potential to significantly improve the performance of PASS.
IV “PASS” Signals for 6G and Beyond
In this section, we will discuss promising applications of PASS to underpin 6G and beyond.
IV-A PASS for Ubiquitous Indoor and Outdoor Coverage
The low-complexity structure of PASS enables it to be easily deployed on diverse structures in both outdoor and indoor scenarios, such as building facades, rooftops, roadside guardrails, and indoor ceilings or walls, as depicted in Fig. 4. PASS address the key challenge of conventional wireless communication, where communication links are often partially or totally blocked by obstacles and suffer from significant long-distance path loss due to those remotely deployed BSs. In particular, PASS establishes LoS-dominated and short-distance wireless channels with the widespread deployment of waveguides and the optimal activation of PAs near the terminals. By preventing wireless signals from being wasted or even interfering other users, PASS is capable of facilitating ubiquitous outdoor and indoor coverage in a cost-efficient and energy-efficient manner. Several attractive applications of PASS in indoor and outdoor scenarios will be discussed below.
On the one hand, PASS is quite appealing in indoor scenarios. As shown in the top right of Fig. 4, PASS installed on the ceiling can flexibly activate PAs to provide wireless communications for users in different rooms. This mitigates the severe penetration losses suffered by wireless signals transmitted by outdoor BSs or indoor access points (APs), particularly at mmWave and THz frequencies. Moreover, by exploiting the near-field channel between PAs and users, efficient beam focusing can be achieved. This not only enhances the desired signal strength but also reduces interference to other users. On the other hand, the widely distributed PASS can also be utilized in outdoor mobile scenarios, such as aerial/space communications and vehicular communications, as illustrated in the middle of Fig. 4. For example, recall the fact that conventional BSs with down-tilt antennas are typically inefficient for aerial communications. As a remedy, PASS installed on rooftops and facades can realize tailored upward beamforming for communications and sensing to effectively support medium- and high-altitude aircrafts as well as low-altitude unmanned aerial vehicles (UAVs) or other high-altitude platforms. As a further advancement, PASS can be employed in space communications. By integrating multiple waveguides and numerous PAs, PASS can form a large-scale antenna array to combat the severe path loss typically encountered in high-frequency satellite communications. Moreover, with PASS uniformly deployed along roadside guardrails, efficient vehicular communications can be achieved by adaptively activating PAs matching with vehicular movement. This is particularly beneficial for autonomous driving in future wireless networks.
Figure 5: Array gain versus the number of pinching antennas. The detailed simulation setup is given in [11].
IV-B Interplay between PASS and NOMA
The combination of PASS and non-orthogonal multiple access (NOMA) provides a “win-win” solution for 6G and beyond. On the one hand, since multiple PAs on a specific waveguide can only be fed with the same set of signals, how to serve multiple users simultaneously becomes a critical challenge. To address this, it is natural to use NOMA to multiplex multiple users’ signal in the power domain over the same radio resource [9]. On the other hand, PASS also facilitates flexible NOMA operations. Recall the fact that NOMA relies on successive interference cancellation (SIC) to handle co-channel interference and therefore SIC order significantly affects the communication performance achieved by each NOMA user. Conventionally, a fixed distance-based SIC order is employed for NOMA, where users closer to the BS, so called strong users, are assigned higher SIC orders due to better channel conditions. In contrast, as shown in the bottom right of Fig. 4, PASS-based NOMA enables a more flexible SIC operation. This is achieved by dynamically adjusting users’ channel conditions through the allocation of PAs and the design of their activation locations. For example, users originally farther from the BS/AP can have higher channel conditions with more numbers of PAs allocated. By doing so, more customized wireless services can be provided by the PASS-based NOMA.
IV-C PASS for NISE/ISAC
In addition to enhancing communication, PASS significantly improves sensing quality and coverage. Since the links between PAs and sensing targets follow the near-field model, PASS enables near-field sensing (NISE) to precisely estimate angular and range information, as well as target velocity [15]. As illustrated in the top left of Fig. 4, two practical structures can be employed: ISAC-PASS and NISE-PASS. (1) ISAC-PASS allocates part of the waveguides to transmit communication signals while using the rest to receive echo-sensing signals, enabling simultaneous near-field communication and sensing within the area of interest. (2) NISE-PASS, equipped exclusively with receiving waveguides, operates in a low-power mode to facilitate NISE. By leveraging the widespread deployment of PASS and successively or periodically activating sensing PAs, ubiquitous and highly accurate NISE can be achieved. This capability is particularly beneficial for real-world applications, such as tracking target trajectories or monitoring malicious entities.
Figure 6: Average transmit power versus the minimum SINR.
IV-D PASS for WPT
Wireless power transfer (WPT) is an efficient technology that uses RF signals to supply power to energy-constrained devices and extend their operational lifetime, especially for Internet-of-Things (IoT) and sensor networks. However, conventional WPT is hindered by the challenge of near-far unfairness, where devices located farther from the wireless power-supply BS or AP experience reduced energy-harvesting efficiency due to degraded channel conditions. As shown in the middle right of Fig. 4, this issue can be resolved through PASS-empowered WPT. In particular, tailored PAs can be activated right to each device, which ensures high-quality WPT for all devices, regardless of their locations.
IV-E PASS for PLS
PASS can also enhance secure information transmission in the context of physical layer security (PLS). As shown in the bottom left of Fig. 4, PASS enables legitimate information transmission to effectively bypass the eavesdropper. On the one hand, since signal transmission predominantly occurs within the waveguide of PASS, the likelihood of legitimate information being eavesdropped is significantly reduced compared to broadcast wireless channels. On the other hand, for the short-distance PA-user transmission, the near-field beam focusing can be employed to minimize the information leakage.
V Numerical Case Studies
In this section, we present numerical examples to characterize the performance of PASS employing pinching beamforming and to compare them with the conventional baselines.
V-A Array Gains of PASS
The array gain of PASS is analyzed by considering a setup where multiple PAs, equally spaced along a single waveguide, are used to serve a single-antenna user; see [11, Figure 1]. This setup represents the most fundamental pinching beamforming scenario, and the corresponding results serve as a lower bound for those more advanced designs.
Unlike conventional fixed-location antenna systems, the wireless propagation in PASS is LoS-dominated and near-field-driven. Considering these factors, the array gain of the system has been analyzed with respect to the number of PAs, as illustrated in Figure 5 [11]. The findings in Figure 5 indicate that the array gain does not increase monotonically with the number of antennas. Instead, there exists an optimal number of PAs that maximizes the array gain achievable by PASS. Notably, the maximum array gain exceeds that of conventional-antenna systems. Since the considered naive pinching beamforming represents a lower bound for other potential designs, these observations lead to the following conclusions: i) Due to their flexibility, PASS can achieve a higher array gain than conventional fixed-location antenna systems. ii) Optimizing the number of PAs is essential for PASS to fully realize their performance potential.
V-B Transmit Power Minimization by PASS
Although PASS has demonstrated the ability to achieve enhanced array gains, the design of the joint transmit and pinching beamforming for maximizing the PASS performance is a non-trivial task. The key challenges stem not only from the coupling between transmit and pinching beamforming but also from the combined impact of all PAs on both the amplitude and phase of the signals. To tackle these challenges, we propose leveraging penalty dual decomposition (PDD) to effectively decompose the optimization parameters.
To evaluate the effectiveness of PDD in enhancing the performance of multi-user PASS, we examine a transmit power minimization problem constrained by the minimum signal-to-interference-plus-noise ratio (SINR) requirement. We assume a 15 GHz system where five 50-meter waveguides are deployed in parallel with a spacing of 5 meters, serving four users randomly distributed 15-20 meters away from the BS. Each waveguide has 6 PAs, and a sub-connected multiplexing transmission architecture is used. For comparison, we consider two baseline systems: fully-digital MIMO with 5 antennas and massive MIMO with 30 antennas and 5 RF chains. The numerical results obtained by the Monte Carlo method are illustrated in Fig. 6, where the number of available PA positions for discrete activation is set to 200. These results highlight the significant performance gains of PASS in reducing transmit power, even when compared to massive MIMO with a more complex transmission architecture. Additionally, it is interesting to observe that the performance loss caused by the proportional power model is negligible compared to the equal power model.
VI Conclusions and Open challenges
In this article, the promising concept of PASS was introduced, highlighting their great potential for enabling efficient wireless communications over the “last meter”. The basic physical principle, signal model, and communication design of PASS have been discussed. By leveraging the novel pinching beamforming, two practical transmission architectures were developed for PASS, whose advantages and limitations were also described. Moreover, a range of promising applications for employing PASS in wireless networks were explored. Numerical case studies demonstrated the superiority of PASS over conventional multi-antenna transmission. As PASS remains in its early stages of research, there are numerous open challenges and opportunities to unlock its full potential. Several of these key research directions are exemplified below:
  • Spatial Modeling for PASS with Stochastic Geometry: Stochastic geometry provides a powerful framework for evaluating the average performance of wireless networks, which can be used to determine key design metrics, such as the required number of PAs, the ratio between PAs and users, and optimal placement strategies. However, the unique features of PASS—such as near-field effects, LoS-dominated links, and location-flexible pinching—necessitate the development of more accurate spatial models to capture the spatial randomness, which can be a promising and interesting research direction.
  • Machine Learning Empowered PASS: Machine learning empowers both channel estimation and fast pinching deployment for PASS. Owing to the large spatial extent, PASS may observe non-stationary multi-path characteristics in different serving areas. Generative learning can capture the spatial non-stationarity and reconstruct large-scale PASS channels. Furthermore, low-complexity learning-based designs are required for user-centric pinching of a large number of PAs, thus manipulating PAs for both path loss compensation and constructive interference.
  • Advanced Hardware Designs for PASS: Pinching beamforming plays a critical role in determining the flexibility of PASS and the maximum achievable performance gain. Advanced PASS hardware designs are essential in this context, posing several challenges, such as real-time and continuous control of PA movements or activation, active control of power leakage from PAs, mitigation of unwanted power losses, and desired dielectric materials. This requires multi-disciplinary research efforts.
Biographies
Yuanwei Liu (Fellow, IEEE) is a (tenured) Full Professor at The University of Hong Kong.
Zhaolin Wang (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Xidong Mu (Member, IEEE) is a Lecturer with the Centre for Wireless Innovation (CWI), Queen’s University Belfast.
Chongjun Ouyang (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Xiaoxia Xu (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Zhiguo Ding (Fellow, IEEE) is a Professor in Communications at the University of Manchester and a Distinguished Adjunct Professor at Khalifa University.